In a​ one-tail hypothesis test where you reject Upper H 0 only in the upper ​tail, what is the p dash value if Upper Z Subscript STAT Baseline equals plus 2.40 question markIn a one-tail hypothesis test where you reject H, only in the upper tail, what is the p-value if ZSTAT = + 2.40? Click here t
Answer #1

solution

Z= +2.40

P(z >2.40 ) = 1 - P(z <2.40 ) = 1 -0.9918=0.0082

p-value=0.0082 ( using standard normal table )